QUESTION 24

  1. In 1980 the price of a microwave oven was $450 and CPI was 82. Today the price of a microwave overn is $100 and CPI is 259.

    Round to two decimal places.

    What is the 1980 price of a microwave oven measured in todays dollars? $

    What is todays price of a microwave oven measured in 1980 dollars? $

4 points   

QUESTION 25

  1. The U.S. dollar is called

    faith money.

    fiat money because the law decrees it is money.

    convertible money because the government stands ready to convert it into gold or silver.

    commodity money, because it is convertible into gold.

2 points   

QUESTION 26

  1. M1 is composed of

    currency inside of banks, traveler's checks, and government-issued checks.

    traveler's checks, credit cards, and e-cash.

    currency held by individuals and businesses, traveler's checks, and checkable deposits owned by individuals and businesses.

    currency held by individuals and businesses, traveler's checks, and the credit line on credit cards.

2 points   

QUESTION 27

  1. Required reserve ratios are the minimum amount of

    reserves any one bank must hold as a percentage of its deposits.

    deposits any one bank is allowed to accept as percentage of its capital.

    deposits any one bank must hold as a percentage of its reserves.

    reserves any one bank must hold as a percentage of its loans.
Answer #1

Answer

24

CPI increased by a factor of (259-82)/80 = 2.2125

Price of 1980 microwave today 82 * 2.2125 = 181.42

CPI Decrease from today to 1980 (259-82)/259 = 0.6833

todays price of a microwave oven measured in 1980 dollars = 100 * 0.6833 = 68.33

(First question and its subparts solved as HOMEWORKLIB's policy.Please re-upload the rest of the questions for answering. Consider giving an upvote if you find it useful)